An artificial satellite of earth is revolving round earth at a height 900000m from the earth surface. Find the minimum velocity of the satellite & period of revolution.


1
Expert's answer
2021-06-28T17:11:04-0400
"v=\sqrt{\frac{GM}{R+h}}\\v=\sqrt{\frac{(6.67\cdot10^{-11})(6\cdot10^{24})}{6.38\cdot10^6+9\cdot10^5}}=7414\frac{m}{s}"

"T=\frac{2\pi (6.38\cdot10^6+9\cdot10^5)}{7414}=6170\ s"